OpenStudy (erinweeks):

What is the solution of each proportion? 4/9=m/54? Anyone know how to set up the solving process?

OpenStudy (erinweeks):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

4/9=m/54 4*54 = 9*m ... cross multiply I'll let you finish

OpenStudy (erinweeks):

216=9m m=24

OpenStudy (erinweeks):

:) Right?

jimthompson5910 (jim_thompson5910):

you got it nice work

OpenStudy (erinweeks):

Thanks Jim is there anyway you could help me with one last question?

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (erinweeks):

 the quetion is are the two triangles similar? How do you know? a.Yes, by SAS~ b. Yes, by SSS~ c. The triangles are not similar d.Yes, by AA~ I just want to make sure im solving this right sorry lol